OpenStudy (anonymous):

Convert decimal to fraction 1. 0.41= 2. 0.158= 3. 0.277= 4. 0.55= 5. 0.57= 6. 0.198=

OpenStudy (anonymous):

1. 0.41= 41/100 2. 0.158= 158/1000=79/500 3. 0.277= 277/1000 4. 0.55= 55/100=11/20 5. 0.57= 57/100 6. 0.198=198/1000=99/500

OpenStudy (goformit100):

http://www.webmath.com/

OpenStudy (anonymous):

Thx guise :)